Using the colored bands on the two resistors at your table, you determine that the two resistances of the resistors are 100 ohms with 20% tolerance and 330 ohms with a 10% tolerance. What would you expect the equivalent resistance of the two resistors to be (including the uncertainty) if they are connected in parallel?
Answer #1

R1 = 100 +/- 20 ohm

R2 = 330 +/- 33 ohm

Parallel resistence= (R1xR2)/(R1+R2)

Parallel resistence= (100 * 330) /(100+330) = 76.74


Parallel resist. error= ( R12 dB + R22 dA ) / (R1+ R2)2
where R1=100.0 +/- 20 ohm and R2=330.0 +/- 33 ohm

error=(1002 x 20+3302 x 33)/(100 + 330)2= 20.517 ohm
therefore Reqv = 76.74 +/- 20.517 ohm